# Correlated Noise in Epoch-Based Stochastic Gradient Descent: Implications for Weight Variances

### v2

@@ 4.3 Variance for late training phase
Relations for large Hessian eigenvalues: For Hessian eigenvectors with eigenvalues $\lambda_{i}>\lambda_{\textrm{cross}}$ and when $M\gg 1/(1-\beta)$, the effects of noise anti-correlations are minimal. Consequently, we can use the following approximate relationships, which also hold true in the absence of correlations:
$\sigma_{\theta,i}^{2}$ $\approx\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{(1-\beta)\bigl{(}2(1+\beta)-\eta\lambda_{i}\bigr{)}}\cdot\frac{1+\beta}{\eta\lambda_{i}}\ ,$ (11a)
$\sigma_{v,i}^{2}$ $\approx\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{(1-\beta)\bigl{(}2(1+\beta)-\eta\lambda_{i}\bigr{)}}\cdot 2\ ,$ (11b)
$\tau_{i}$ $\approx\frac{1+\beta}{\eta\lambda_{i}}\ .$ (11c)
The detailed derivation of these formulas is presented in Section C.3.
It is frequently the case that the product $\eta\lambda_{i}$ is considerably less than one, enabling us to further simplify the prefactor of the variances. Assuming the noise covariance matrix is proportional to the Hessian matrix, such that $\sigma_{\delta g,i}^{2}\propto\lambda_{i}$, we derive the following power laws for the variances: $\sigma_{\theta,i}^{2}\propto\text{const.}$. and $\sigma_{v,i}^{2}\propto\lambda_{i}$. As a result, in the subspace spanned by Hessian eigenvectors with eigenvalues $\lambda_{i}>\lambda_{\textrm{cross}}$, our theory predicts an isotropic weight variance $\bm{\mathbf{\Sigma}}$.
@@
$\sigma_{\theta,i}^{2}$ $\approx\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{2(1-\beta)(1+\beta)}\cdot\frac{M}{3}\frac{1+\beta}{1-\beta}\ ,$ (12a)
$\sigma_{v,i}^{2}$ $\approx\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{2(1-\beta)(1+\beta)}\cdot 2\ ,$ (12b)
$\tau_{i}$ $\approx\frac{M}{3}\frac{1+\beta}{1-\beta}\ .$ (12c)
The derivation of these formulas is provided in Section C.3.
(a) Variances
(b) Correlation time
Figure 2: Relationship between Hessian eigenvalues and the variances of weights and velocities, as well as correlation times. The mean velocity of the weight trajectory has been subtracted (see Section 5.3). In the upper panel, we present the variances of weights and velocities. The solid lines signify the regions utilized for a linear fit. The exponents resulting from the power law relationship are $1.077\pm 0.012$ for weight variance and $1.066\pm 0.002$ for velocity variance, with a $2\sigma$-error. Our theory suggests these exponents should be equal to one. The lower panel showcases the correlation time together with the theoretical prediction resulting from Equation 9.
@@
If we once again assume that the noise covariance matrix is proportional to the Hessian matrix, such that $\sigma_{\delta g,i}^{2}\propto\lambda_{i}$, we obtain the following power laws for the variances: $\sigma_{\theta,i}^{2}\propto\lambda_{i}$ and $\sigma_{v,i}^{2}\propto\lambda_{i}$. This implies that in the subspace spanned by Hessian eigenvectors with eigenvalues $\lambda_{i}<\lambda_{\textrm{cross}}$, the weight variance $\bm{\mathbf{\Sigma}}$ is not isotropic but proportional to the Hessian matrix $\bm{\mathbf{H}}$.
It is s noteworthy that the correlation time $\tau_{i}\approx\frac{M}{3}\frac{1+\beta}{1-\beta}\eqqcolon\tau_{\textrm{SGD}}$ in this subspace is independent of the Hessian eigenvalue. Moreover, $\tau_{\textrm{SGD}}$ is equivalent to the correlation time of the noise $M/3$, up to a factor that depends on momentum.
@@ 5.1 Analysis Setup
In order to corroborate our theoretical findings, we have conducted a small-scale experiment. We have trained a LeNet architecture, similar to the one described in [32], using the Cifar10 dataset [35]. LeNet is a compact convolutional network comprised of two convolutional layers followed by three dense layers. The network comprises approximately 137,000 parameters. As our loss function, we employed Cross Entropy, along with an L2 regularization with a prefactor of $10^{-4}$. Here we present results for a single seed and specific hyperparameters. However, we have also performed tests with different seeds and combinations of hyperparameters, all of which showed comparable qualitative behavior (see Appendix J). Furthermore, in Appendix L we studied a ResNet architecture [5], a different and more modern network, where we obtained similar results.
We used SGD to train the network for 100 epochs, employing an exponential learning rate schedule that reduces the learning rate by a factor of $0.98$ each epoch. The initial learning rate is set at $5\cdot 10^{-3}$, which eventually reduces to approximately $7\cdot 10^{-4}$ after 100 epochs. The momentum parameter and the minibatch size $S$ are set to $0.9$ and $50$, respectively, which results in a thousand minibatches per epoch, $M=1000$. This setup achieves 100% training accuracy and 63% testing accuracy. The evolution of loss and accuracy during training can be seen in Appendix I. We then compute the variances right after the initial schedule over a period of 20 additional epochs, equivalent to 20,000 update steps. Throughout this analysis period, the learning rate is maintained at $7\cdot 10^{-4}$ and the recorded weights are designated by $\vec{\theta}_{k}$, with $k=K,\dots,K+T$ and $T=20{,}000$.
Given the impracticability of obtaining the full covariance matrix for all weights and biases over this period due to the excessive memory requirements, we limit our analysis to a specific subspace. We approximate the five thousand largest eigenvalues and their associated eigenvectors of the Hessian matrix $\vec{H}(\vec{\theta}_{K})$, drawn from the roughly 137,000 total, using the Lanczos algorithm [36]. Here, $\vec{\theta}_{K}$ represents the weights at the beginning of the analysis period. To compute the Hessian-vector products required for the Lanczos algorithm, we employ the resource-efficient Pearlmutter trick [37]. The eigenvectors of the Hessian matrix are represented by $\bm{\mathbf{p_{i}}}$, and the projected weights by $\theta_{k,i}=\vec{\theta}_{k}\cdot\vec{p_{i}}$. The variances are computed exclusively for these particular directions. The distribution of the approximated 5,000 eigenvalues is illustrated in Appendix G.
@@ 5.2 Noise Autocorrelations
We scrutinize the correlations of noise by recording both the minibatch gradient $\vec{g}_{k}(\vec{\theta}_{k})$ and the total gradient $\vec{\nabla}L(\vec{\theta}_{k})$ at each update step throughout the analysis period, enabling us to capture the actual noise term $\bm{\mathbf{\delta g}}_{k}(\vec{\theta}_{k})$. All these are projected onto the approximated Hessian eigenvectors.
The theoretical prediction for the anti-correlation of the noise is proportional to the inverse of the number of batches per epoch, in our case on the order of $10^{-3}$. To extract the predicted relationship from the fluctuating data, we compute the autocorrelation of the noise term for each individual Hessian eigenvector. We then proceed to average these results across the 5,000 approximated eigenvectors. Figure 1 provides a visual representation of this analysis, showcasing a strong alignment between the empirical autocorrelation of noise and the prediction derived from our theory. This consistency can be interpreted as a validation of our assumption that the noise is spatially independent.
@@ 5.3 Variances and Correlation time
Previous studies have observed that network weights continue to traverse the parameter space even after the loss appears to have stabilized [19, 28, 32]. This behavior persists despite the use of L2 regularization and implies that the recorded weights, $\vec{\theta}_{k}$, do not settle into a stationary distribution. Notably, however, over the course of the 20 epochs under scrutiny, the weight movement, excluding the SGD noise, appears to be approximately linear in time. This suggests that the mean velocity $\bm{\mathbf{\bar{v}}}\coloneqq\langle\bm{\mathbf{v}}_{k}\rangle$ is substantial compared to the SGD noise. To isolate this ongoing movement and uncover the underlying structure, we redefine $\vec{\theta}_{k}$ and $\vec{v}_{k}$ by subtracting the mean velocity. This results in $\vec{\theta}_{k}^{\textrm{(s)}}\coloneqq\vec{\theta}_{k}-\vec{\bar{v}}\cdot k$ and $\vec{v}_{k}^{\textrm{(s)}}\coloneqq\vec{v}_{k}-\vec{\bar{v}}$. We then compute the variances of these redefined values, $\vec{\theta}_{k}^{\textrm{(s)}}$ and $\vec{v}_{k}^{\textrm{(s)}}$, which exhibit a more stationary distribution.
Again, we limit our variance calculations to the directions of the 5,000 approximated Hessian eigenvectors. In the two different regimes of Hessian eigenvalues, either greater or lesser than the crossover value $\lambda_{\textrm{cross}}$, the weight and velocity variance closely follow the respective power law predictions from our theory (see upper panel of Figure 2). The slight discrepancy, where the predicted exponent of one does not lie within the error bars, may arise from minor deviations in the noise covariance from the Hessian approximation $\vec{C}\propto\vec{H}$. The calculated correlation time, derived from the ratio between the weight and velocity variance, aligns reasonably well with our theoretical predictions (see lower panel of Figure 2). This correlation time prediction remains independent of the exact relation between $\vec{C}$ and $\vec{H}$, thereby providing a more general result.
6 Limitations
As highlighted in Section 5.3, the weight distribution is not entirely stationary; the weights exhibit a finite mean drift. To fully understand the underlying weight variances, it is essential to account for this mean velocity. However, for time windows extending beyond the measured 20 epochs, the mean velocity’s variability increases. This change can cause the average weight movement to deviate from linearity, potentially leading to higher variances in flat directions than predicted.
In general, the late phase of training under scrutiny may not be representative of the entire training process. However, we argue that exploring such a regime, especially where a quadratic approximation of the loss seems feasible, can provide important insights into the dynamics of SGD, and in particular its stability around a minimum. Furthermore, it is an important starting point for our research, since such a situation can be treated analytically.
@@ 7 Discussion
We provide an intuitive interpretation of our results through the correlation time associated with different Hessian eigenvectors and corroborate this timescale empirically. A natural question then emerges about how this correlation time relates to the total duration of the analysis period. Indeed, the maximum correlation time, $\tau_{\textrm{SGD}}$, predicted for the flat directions is roughly seven epochs, comparable to the recorded 20-epoch period. However, the outcomes differ considerably when we sample the examples without replacement during training (see Appendix H). Hence, we infer that the observed correlation time behavior stems from the correlations resulting from sampling without replacement.
@@
Here, we primarily focus on the weight covariance denoted as $\bf\Sigma$, treating the correlation time $\tau_{i}$ and the velocity variance $\bf\Sigma_{v}$ as auxiliary variables. Our assumption is that the velocity covariance $\bf\Sigma_{v}$ is directly proportional to the noise covariance $\bf C$. However, the behavior of the weight covariance $\bf\Sigma$ is more complex. This complexity arises from the autocorrelation time, which links the two variances. Specifically, in the context of a single eigendirection, the relationship is expressed as $\sigma_{\theta,i}^{2}=\frac{1}{2}\tau_{i}\sigma_{v,i}^{2}$. A detailed examination of these individual quantities offers further insights.
Additionally, our findings for the three quantities contrast with a previous empirical study [32]. In Appendix F, we clarify how the different analysis method used in that study impacts the results due to finite size effects.
@@
In this manuscript, we study the influence of smaller Hessian eigenvalues and corresponding eigendirections, which are often overlooked in discussions about the characteristics of a minimum in optimization landscapes. Previous research primarily focuses on the role of larger Hessian eigenvalues in defining crucial properties of a minimum [38, 39]. Our work raises the question of the potential impact of the lower-than-expected weight variance in flat directions, which we argue are significant in this scenarios. We propose that this reduced variance might enhance the stability of a minimum, particularly given the substantial difference in the number of flat and steep directions, as illustrated in Appendix G. This suggests that the observed reduction in variance helps explain why SGD exploration predominantly focuses on a limited subspace of Hessian eigenvectors with higher curvature, as reported in prior studies [38, 18].
To further substantiate this argument, we examine the impact of weight variance on the additional training error. In a simplified scenario with a quadratic loss for a Hessian eigenvector direction with eigenvalue $\lambda_{i}$ and Gaussian weight fluctuations of variance $\sigma_{\theta,i}^{2}$, the expected loss variance due to fluctuations is $l_{{\rm fluct},i}=\frac{1}{2}\lambda_{i}\sigma_{\theta,i}^{2}$. Although the contribution of flat directions to the total fluctuation of the loss $l_{\rm fluct}=\sum_{i}l_{{\rm fluct},i}$ is reduced by their small Hessian eigenvalues, this is counterbalanced by their abundance. Analyzing the 5,000 largest Hessian eigenvalues of the LeNet, we find that our variance prediction leads to a 62% reduction in loss fluctuation compared to the constant weight variance assumption, despite only affecting the smaller Hessian eigenvalues. We note that this effect would only be stronger if we also considered the remaining 132,000 Hessian eigendirections.
Furthermore, we argue that flat directions are significant for generalization, a point elaborated in Appendix K. Our investigation of LeNet’s test accuracy, following the decomposition of neural network weights in the Hessian eigenbasis, reveals compelling insights. Omitting the projection of the weight vector onto the top 35 eigenvectors only reduces test accuracy from 63% to 54%. Excluding the top 5,000 eigenvectors, as identified in our analysis, leads to a further drop to 44%. The accuracy only falls to 10%, akin to random guessing, after discarding around 7,000 eigenvectors. These findings underscore the relevance of small Hessian eigenvalues and weight variances for controlling fluctuations in the magnitude of weights, thereby affirming the importance of weight fluctuations in these flat directions.

### v3

@@ 4.3 Variance for late training phase
Relations for large Hessian eigenvalues: For Hessian eigenvectors with eigenvalues $\lambda_{i}>\lambda_{\textrm{cross}}$ and when $M\gg 1/(1-\beta)$, the effects of noise anti-correlations are minimal. Consequently, we can use the following approximate relationships, which also hold true in the absence of correlations and, therefore, also when drawing examples with replacement:
$\sigma_{\theta,i}^{2}\approx\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{(1-\beta)\bigl(2(1+\beta)-\eta\lambda_{i}\bigr)}\cdot\frac{1+\beta}{\eta\lambda_{i}}\ ,\ \sigma_{v,i}^{2}\approx\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{(1-\beta)\bigl(2(1+\beta)-\eta\lambda_{i}\bigr)}\cdot 2\ ,$ (8)
and $\tau_{i}\approx\frac{1+\beta}{\eta\lambda_{i}}$ . The detailed derivation of these formulas is presented in Section˜B.3.
It is frequently the case that the product $\eta\lambda_{i}$ is considerably less than one, enabling us to further simplify the prefactor of the variances. Assuming the noise covariance matrix is proportional to the Hessian matrix, such that $\sigma_{\delta g,i}^{2}\propto\lambda_{i}$, we derive the following power laws for the variances: $\sigma_{\theta,i}^{2}\propto\text{constant}$ and $\sigma_{v,i}^{2}\propto\lambda_{i}$. As a result, in the subspace spanned by Hessian eigenvectors with eigenvalues $\lambda_{i}>\lambda_{\textrm{cross}}$, our theory predicts an isotropic weight variance $\boldsymbol{\mathbf{\Sigma}}$.
@@
$\sigma_{\theta,i}^{2}\approx\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{2(1-\beta)(1+\beta)}\cdot\frac{M}{3}\frac{1+\beta}{1-\beta}\ ,\ \sigma_{v,i}^{2}\approx\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{2(1-\beta)(1+\beta)}\cdot 2$ (9)
and $\tau_{i}\approx\frac{M}{3}\frac{1+\beta}{1-\beta}\eqqcolon\tau_{\textrm{SGD}}$ . Therefore, the weight variance is reduced by a factor of $\frac{M\eta\lambda_{i}}{3(1-\beta)}$ compared to the case without anti-correlations. The derivation of these formulas is provided in Section˜B.3.
@@
If we once again assume that the noise covariance matrix is proportional to the Hessian matrix, such that $\sigma_{\delta g,i}^{2}\propto\lambda_{i}$, we obtain the following power laws for the variances: $\sigma_{\theta,i}^{2}\propto\lambda_{i}$ and $\sigma_{v,i}^{2}\propto\lambda_{i}$. This implies that in the subspace spanned by Hessian eigenvectors with eigenvalues $\lambda_{i}<\lambda_{\textrm{cross}}$, the weight variance $\boldsymbol{\mathbf{\Sigma}}$ is not isotropic but proportional to the Hessian matrix $\boldsymbol{\mathbf{H}}$. It is noteworthy that the correlation time $\tau_{i}\approx\frac{M}{3}\frac{1+\beta}{1-\beta}\eqqcolon\tau_{\textrm{SGD}}$ in this subspace is independent of the Hessian eigenvalue. Moreover, $\tau_{\textrm{SGD}}$ is equivalent to the correlation time of the noise $M/3$, up to a factor that depends on momentum. Furthermore, with both approximations in place, one can understand the small-eigenvalue regime as the range of $\lambda$ for which the correlation time induced by the loss landscape exceeds the timescale of the anti-correlations – both scaled appropriately by the momentum factor –i.e., when $\frac{1+\beta}{\eta\lambda}>\frac{M}{3}\frac{1+\beta}{1-\beta}$ for $\lambda<\lambda_{\rm cross}$.
weight variance $\sigma_{\theta,i}^{2}$ velocity variance $\sigma_{v,i}^{2}$ correlation time $\tau_{i}$
large eigenvalues:
$\lambda_{i}>\lambda_{\rm cross}$
$\frac{1}{2}\cdot\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{(1-\beta)(1+\beta)}\cdot\frac{1+\beta}{\eta\lambda_{i}}$ $\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{(1-\beta)(1+\beta)}$ $\frac{1+\beta}{\eta\lambda_{i}}$
small eigenvalues:
$\lambda_{i}<\lambda_{\rm cross}$
$\frac{1}{2}\cdot\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{(1-\beta)(1+\beta)}\cdot\frac{M}{3}\frac{1+\beta}{1-\beta}$ $\frac{\eta^{2}\sigma_{\delta g,i}^{2}}{(1-\beta)(1+\beta)}$ $\frac{M}{3}\frac{1+\beta}{1-\beta}\coloneqq\tau_{\rm SGD}$
Table 1: Summary of variances and correlation time along Hessian eigenvectors, depending on the eigenvalue magnitude. For simplicity, we assume $\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{C}}=\boldsymbol{\mathbf{C}}\boldsymbol{\mathbf{H}}$, $\eta\lambda_{i}\ll 1$ for all directions, and $M\gg 1/(1-\beta)$ with $M=N/S$ the number of minibatches per epoch. The results for large eigenvalues also coincide with those one obtains in the absence of anti-correlations in the noise. Note: fluctuations of the weight vector during an epoch can modify the noise anti-correlations and induce a minimal weight variance at very small eigenvalues, as discussed in ˜E.
We summarize the results of Corollary 4.4 and Corollary 4.5 under simplifying assumptions in Table˜1. Notice that modifications to the noise anti-correlations caused by a fluctuating weight vector can affect the predictions above, particularly for very small Hessian eigenvalues. In ˜E, we show that these modifications introduce a minimal weight variance, below which the variance in very low-curvature directions cannot drop, even as the Hessian eigenvalues approach zero. However, in all our numerical experiments, the Hessian eigenvalues are sufficiently large that this lower-bound regime plays only a negligible role. In particular, in ˜E we derive that for the observed distribution of Hessian eigenvalues, the weight variance is expected to plateau only for eigenvalues $\lambda_{i}\lesssim 10^{-4}\lambda_{\rm max}$, which corresponds to values on the order of $10^{-1}$. In Figure˜2 we observe slight deviations from the predicted variance–curvature proportionality for eigenvalues on the order of $10^{-1}$.
@@ 5.1 Analysis Setup
In order to corroborate our theoretical findings, we have conducted a small-scale experiment. We have trained a LeNet architecture, similar to the one described in [32], using the CIFAR10 dataset [36]. LeNet is a compact convolutional network comprised of two convolutional layers followed by three dense layers. The network comprises approximately 137,000 parameters. Here we present results for a single seed and specific hyperparameters. However, we have also performed tests with different seeds and combinations of hyperparameters, all of which showed comparable qualitative behavior (see Section˜G.5). Furthermore, in Section˜G.8 we studied a ResNet architecture [5], a more modern network, where we obtained similar results.
The training parameters together with the schedule are described in Section˜G.1 and result in a thousand minibatches per epoch, M=1, [^0] M=1{,}000. The setup achieves 95% training accuracy and about 70% testing accuracy. We compute the variances right after the initial schedule over a period of 20 additional epochs, equivalent to 20,000 update steps. During the initial schedule, we use an exponential learning rate decay applied once per epoch, while our theoretical results assume a constant learning rate. To better align the analysis with theory, we maintain a constant learning rate throughout the 20-epoch analysis period. However, since the learning rate changes only slowly across epochs – and remains fixed within each epoch – we expect deviations from theory due to this potential mismatch to be minimal when analyzing timeframes spanning only a few epochs.
The recorded weights are designated by $\vec{\theta}_{k}$, with $k=K,\dots,K+T$ and T=20, [^0] T=20{,}000. Given the impracticability of obtaining the full covariance matrix for all weights over this period due to the excessive memory requirements, we limit our analysis to a specific subspace. Employing the resource-efficient Pearlmutter trick [37], we approximate the 5,000 largest eigenvalues and their associated eigenvectors of the Hessian matrix $\vec{H}(\vec{\theta}_{K})$ at the beginning of the analysis period, drawn from the roughly 137,000 total. The eigenvectors of the Hessian matrix are represented by $\boldsymbol{\mathbf{p_{i}}}$, and the projected weights by $\theta_{k,i}=\vec{\theta}_{k}\cdot\vec{p_{i}}$. The variances are computed exclusively for these particular directions. The distribution of the approximated 5,000 eigenvalues is illustrated in Section˜G.1.
@@ 5.2 Noise Autocorrelations
We scrutinize the correlations of noise by recording both the minibatch gradient $\vec{g}_{k}(\vec{\theta}_{k})$ and the total gradient $\vec{\nabla}L(\vec{\theta}_{k})$ at each update step throughout the analysis period, enabling us to capture the actual noise term $\boldsymbol{\mathbf{\delta g}}_{k}(\vec{\theta}_{k})$. All these are projected onto the 5,000 approximated Hessian eigenvectors, yielding scalar components $\delta g_{k,i}=\boldsymbol{\mathbf{\delta g}}_{k}(\vec{\theta}_{k})\cdot\boldsymbol{\mathbf{p}}_{i}$. The theoretical prediction for the anti-correlation of the noise is a V-shaped structure proportional to the inverse of the number of batches per epoch, $1/M$, which, in our case, is on the order of $10^{-3}$. To extract the predicted relationship from the fluctuating data, we compute the autocorrelation for each direction $i$ individually as $\text{corr}_{\delta g,i}(h)=\frac{1}{\sigma_{\delta g,i}^{2}}\cdot\frac{1}{T-|h|}\sum_{k=K}^{T-|h|}\delta g_{k+|h|,i}\,\delta g_{k,i}$.
Due to the limited resolution from finite sampling, the standard deviation of each correlation estimate is around $1/\sqrt{T}\approx 7\cdot 10^{-3}$ – larger than the signal we aim to resolve. Therefore, we average the autocorrelation results across the 5,000 approximated eigenvectors, giving $\text{corr}_{\delta g,\,avg}(h)$ and reducing the expected noise level to approximately $1/\sqrt{5000\cdot T}\approx 10^{-4}$, which enables detection of the predicted structure. Figure˜1 provides a visual representation of this analysis, showcasing a strong alignment between the empirical autocorrelation of noise and the prediction derived from our theory, confirming the V-shaped theoretical autocorrelation.
@@ 5.3 Variances and Correlation time
Again, we limit our variance calculations to the directions of the 5,000 approximated Hessian eigenvectors. In the two different regimes of Hessian eigenvalues, either greater or lesser than the crossover value $\lambda_{\textrm{cross}}$, the weight and velocity variance closely follow the respective predictions from our theory (see left panel of Figure˜2). We observe slight deviations from the predicted variance–curvature proportionality in the very low-curvature directions, where some weight variances are marginally higher than expected. These deviations may result from slow equilibration in very flat directions, as previous studies have observed that network weights continue to traverse the parameter space even after the loss appears to have stabilized [19, 32, 28]. Another explanation could be corrections to the anti-correlations of the noise due to fluctuating weights. A more detailed discussion and extended analysis, can be found in Section˜G.4. The calculated correlation time, derived from the ratio between the weight and velocity variance, also aligns well with our theoretical predictions (see right panel of Figure˜2). This correlation time prediction remains independent of the exact relation between $\vec{C}$ and $\vec{H}$, thereby providing a more general result.
@@ 6 Discussion
We provide an intuitive interpretation of our results by examining the correlation times associated with different Hessian eigenvectors, which we also corroborate empirically. Although the 20-epoch analysis period may influence the observed correlation times – since it is comparable to the maximum predicted correlation time $\tau_{\textrm{SGD}}$ – we observe considerable differences when we sample examples without replacement during training (see Section˜G.3). This suggests that the observed correlation time behavior arises from the correlations introduced by sampling without replacement.
@@
Throughout our analysis, we primarily focus on the weight covariance matrix $\mathbf{\Sigma}$, considering the correlation time $\tau_{i}$ and the velocity covariance $\mathbf{\Sigma_{v}}$ as auxiliary variables. Our results indicate that the velocity covariance $\mathbf{\Sigma_{v}}$ is directly proportional to the noise covariance $\mathbf{C}$. However, the behavior of the weight covariance $\mathbf{\Sigma}$ is more intricate due to the autocorrelation time, which connects the two variances. Specifically, for a single eigendirection, this relationship is given by $\sigma_{\theta,i}^{2}=\frac{1}{2}\tau_{i}\sigma_{v,i}^{2}$. Examining these quantities in detail provides further insights. Moreover, our findings for these three quantities differ from those reported in a previous empirical study [32]. In Section˜G.2, we explain how the different analysis methods used in that study led to results impacted by finite-size effects.
@@
In this manuscript, we examine the influence of smaller Hessian eigenvalues, which are often overlooked in discussions of minima characteristics in optimization landscapes. While prior work has focused on larger eigenvalues, we highlight the impact of reduced weight variance in flat directions. This observed reduction may help explain why SGD exploration predominantly occurs within a limited subspace of Hessian eigenvectors with higher curvature, as reported in previous studies [38, 17].
We further connect our theoretical results to the long-standing hypothesis in deep learning that posits that flatter minima tend to generalize better than sharper ones. From this perspective, it is desirable for an optimization algorithm to escape sharp minima more efficiently than flat ones. To assess this behavior, we consider the implications of our result on the escaping efficiency $E_{t}\coloneqq\langle L(\boldsymbol{\mathbf{\theta}}_{t})-L(\boldsymbol{\mathbf{\theta}}_{0})\rangle_{\text{batch noise}}$, a concept introduced in prior work [23, 30], which approximates the expected increase in loss due to stochastic weight fluctuations during training. Under a quadratic approximation of a minimum, $L(\boldsymbol{\mathbf{\theta}})\approx\frac{1}{2}(\boldsymbol{\mathbf{\theta}}-\boldsymbol{\mathbf{\theta}}_{0})^{\rm T}\boldsymbol{\mathbf{H}}(\boldsymbol{\mathbf{\theta}}-\boldsymbol{\mathbf{\theta}}_{0})+L_{0}$, the stationary escaping efficiency becomes $E_{\infty}\approx\frac{1}{2}\mathrm{Tr}[\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{\Sigma}}]$, where $\boldsymbol{\mathbf{\Sigma}}$ is the stationary weight covariance. For clarity, we consider our theoretical predictions for $\boldsymbol{\mathbf{C}}=c_{0}\boldsymbol{\mathbf{H}}$ and without momentum, but similar qualitative behavior holds in the momentum case. When training with SGD with replacement, the weight variance is predicted to be isotropic, $\boldsymbol{\mathbf{\Sigma}}^{\rm w.r.}=\frac{\eta c_{0}}{2}\boldsymbol{\mathbf{1}}$, leading to an escaping efficiency $E_{\infty}^{\rm w.r.}=\frac{\eta c_{0}}{4}\mathrm{Tr}[\boldsymbol{\mathbf{H}}]$. For SGD without replacement, however, the weight variance becomes anisotropic. In directions corresponding to eigenvalues $\lambda_{i}>\lambda_{\rm cross}$, the variance matches that of the with-replacement case, i.e., $\boldsymbol{\mathbf{\Sigma}}_{\text{\tiny(\textgreater)}}^{\rm w/o.r.}=\frac{\eta c_{0}}{2}\boldsymbol{\mathbf{1}}$. In contrast, for eigenvalues $\lambda_{i}<\lambda_{\rm cross}$, the variance becomes $\boldsymbol{\mathbf{\Sigma}}_{\text{\tiny(\textless)}}^{\rm w/o.r.}=\frac{\eta c_{0}}{2}\frac{1}{\lambda_{\rm cross}}\boldsymbol{\mathbf{H}}_{\text{\tiny(\textless)}}$, resulting in a reduced escaping efficiency $E_{\infty}^{\rm w/o.r.}=\frac{\eta c_{0}}{4}\mathrm{Tr}\big[\boldsymbol{\mathbf{H}}_{\text{\tiny(\textgreater)}}\big]+\frac{\eta c_{0}}{4}\frac{1}{\lambda_{\rm cross}}\mathrm{Tr}\big[\boldsymbol{\mathbf{H}}_{\text{\tiny(\textless)}}^{2}\big]$. Since $\boldsymbol{\mathbf{H}}_{\text{\tiny(\textless)}}$ represents the contribution from flatter directions (i.e., those with $\lambda_{i}<\lambda_{\rm cross}$), this result suggests that SGD without replacement is less prone to escaping flat minima than SGD with replacement – particularly given that the majority of Hessian eigenvalues lie below the crossover value in our numerical experiments (see Figure˜5). To demonstrate significance, we analyze the 5,000 largest Hessian eigenvalues of LeNet, include the effects of momentum, and find that our model predicts a 62% reduction in escaping efficiency for the case without replacement compared to sampling with replacement. This effect would be even stronger if all 137,000 directions were considered. This selective suppression of variance in low-curvature directions may help the optimizer remain in flatter regions of the loss landscape, potentially contributing to improved generalization performance. Making this connection between suppressed variance and convergence towards flatter minima more explicit is a direction for future research.
To further assess the impact of anti-correlations in the gradient noise, we investigated the difference in generalization performance between drawing batches in SGD with and without replacement. We trained the network described in Section˜5.1 using the same training schedule and 20 different random seeds, considering the maximum test accuracy computed after each epoch. We found that training without replacement yielded test accuracies that were 0.8% ± 0.2% higher than training with replacement. Specifically, the maximum test accuracy for SGD without replacement was 69.8% on average, compared to 69.0% for SGD with replacement.
We relate this result directly to the anti-correlations we have described by comparing it with a prior study by Orvieto et al. [33]. In that study, the authors considered full-batch gradient descent with artificially added noise that is anti-correlated in time. This noise was found to be beneficial for test accuracy and led to flatter minima. While the anti-correlations they considered have a very short correlation time, they are otherwise very similar to those inherent in SGD without replacement. We therefore propose that the positive effects described in their study could be extended to SGD without replacement due to these anti-correlations. Similarly, Beneventano [39] studies SGD without replacement, showing that it regularizes the trace of the noise covariance along flat directions, leading to a faster escape from saddle points and a better generalization than SGD with replacement. Also, our findings on the impact of anti-correlated gradient noise in SGD without replacement relate to prior work on random reshuffling in convex optimization. In particular, Gürbüzbalaban et al. [40] analyzes convergence rates under strong convexity and shows that without-replacement sampling can accelerate optimization.